Use the five-step method to solve.
One-third of a number is 5 less than half of the same number. What is the number?

Answer by jainenderkapoor(61)


Let the number be = `x`
We have x/3 = x/2 - 5
x/3 - x/2 = -5
(2x - 3x)/6 = -5
-x = -5*6
-x = -30

x = 30


I hope that the solution is clear to you.
If you have any doubt you are welcome to contact me.
I am also available for online tutoring in math at very reasonable rates.
My mail id is jainenderkapoor@gmail.com
Jainender Kapoor